OpenStudy (wow.):

Marneshia walked 3/8 of a mile in of 3/5 an hour. What equation can be used to calculate her unit rate in miles per hour? (5 points) 3/5 3/8 =1 3/5 miles per hour 3/8 3/5=5/8 miles per hour 3/8 3/5=1 3/5 miles per hour 3/5 3/8= 5/8mile per hour

Vocaloid (vocaloid):

hint: divide miles/hours

OpenStudy (anonymous):

D. 3/5 3/8= 5/8mile per hour

OpenStudy (anonymous):

3/8 as a decimal is 0.375 3/5 as a decimal is 0.6 0.375/0.6=0.625 D. how it has 5/8, turn that into a decimal 5/8 as a decimal is 0.625 its equal to 0.375/0.6=0.625 your answer is D. D. 3/5 3/8= 5/8mile per hour